# Mechanical quadrature squeezing beyond the 3dB limit via quantum learning control

### v1

@@ 4 Discussions and conclusion
For generation of mechanical quadrature squeezing, we need to minimize the loss function $\Delta X_{b}^{2}(\theta,T)$. In this Appendix, we will expound in detail how to achieve this goal using the gradient descent algorithm. The mathematical expression of the gradient descent algorithm for mechanical squeezing generation is
@@
According to Eq. (17), we can further calculate the results of $\delta\mathrm{V}_{33}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{44}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{34}(T)/\delta\mathcal{Q}(s)$, and $\delta\mathrm{V}_{43}(T)/\delta\mathcal{Q}(s)$. The variation of the covariance-matrix elements with respect to the pulse amplitude $\Omega(t)$ and phase $\phi(t)$ can be expressed as a linear combination of the variation of these second-order moments with respect to driving amplitude $\Omega(t)$ and phase $\phi(t)$,
@@
It can be seen from Eq. (18) that, to obtain the result in Eq. (17), we need to further calculate the variations of these second-order moments with respect $\Omega(t)$ and $\phi(t)$. This can be achieved by taking the variation with respect to $\Omega(s)$ and $\phi(s)$ on both sides of Eq. (8), namely,
@@
Since $\mathcal{Q}(s)$ is a function of $s$, $\delta\dot{\mathbf{X}}(t)/\delta\mathcal{Q}(s)$ in Eq. (19) can be expressed as $\frac{d}{dt}\frac{\delta{\mathbf{X}}(t)}{\delta\mathcal{Q}(s)}$ [72]. Under the initial conditions $\delta\mathbf{X}(0)/\delta\mathcal{Q}(s)=[0,...,0]_{10\times 1}^{\text{T}}$, the solution of Eq. (19) at the target time $T$ can be obtained as
@@
To know the explicit expressions for $\delta\mathbf{M}(s)/\delta\mathcal{Q}(s)$ and $\delta\mathbf{N}(s)/\delta\mathcal{Q}(s)$ in Eq. (Appendix. Derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ in gradient descent algorithm), we need to obtain the values of $\delta\alpha(s)/\delta\mathcal{Q}(s)$, $\delta\beta(s)/\delta\mathcal{Q}(s)$, and their Hermitian conjugate. Taking the variation with respect to $\mathcal{Q}(s)$ on both sides of Eqs. (7a) and (7b) as well as their Hermitian conjugate equations, we have
@@
$\mathbf{A}(t)=\left[\frac{\delta\alpha(t)}{\delta\mathcal{Q}(s)},\frac{\delta\beta(t)}{\delta\mathcal{Q}(s)},\frac{\delta\alpha^{\ast}(t)}{\delta\mathcal{Q}(s)},\frac{\delta\beta^{\ast}(t)}{\delta\mathcal{Q}(s)}\right]^{\text{T}},$ (22)
@@
$\left[\frac{\delta\alpha(s)}{\delta\Omega(s)},\frac{\delta\beta(s)}{\delta\Omega(s)},\frac{\delta\alpha^{\ast}(s)}{\delta\Omega(s)},\frac{\delta\beta^{\ast}(s)}{\delta\Omega(s)}\right]^{\text{T}}=\left[\frac{1}{2}ie^{-i\phi(s)},0,-\frac{1}{2}ie^{i\phi(s)},0\right]^{\text{T}},$ (26a)
$\left[\frac{\delta\alpha(s)}{\delta\phi(s)},\frac{\delta\beta(s)}{\delta\phi(s)},\frac{\delta\alpha^{\ast}(s)}{\delta\phi(s)},\frac{\delta\beta^{\ast}(s)}{\delta\phi(s)}\right]^{\text{T}}=\left[\frac{1}{2}\Omega(s)e^{-i\phi(s)},0,\frac{1}{2}\Omega(s)e^{i\phi(s)},0\right]^{\text{T}}.$ (26b)
Thus, we obtain the variation of the displacement amplitudes $\alpha(s)$ and $\beta(s)$ with respect to the driving amplitude $\Omega(s)$ and phase $\phi(s)$, respectively. Based on the value of $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}(s)$, we can further perform the gradient descent algorithm until a satisfactory value of the loss function $\Delta X_{b}^{2}(\theta,T)$ is obtained.

### v2

@@ Appendix A Derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$
For generation of mechanical quadrature squeezing, we need to minimize the loss function $\Delta X_{b}^{2}(\theta,T)$. In this Appendix, we will clarify in detail how to achieve this goal using the gradient-descent algorithm. The mathematical expression of the gradient-descent algorithm for mechanical squeezing generation is
@@
According to Eq. (17), we can further calculate the results of $\delta\mathrm{V}_{33}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{44}(T)/\delta\mathcal{Q}(s)$, $\delta\mathrm{V}_{34}(T)/\delta\mathcal{Q}(s)$, and $\delta\mathrm{V}_{43}(T)/\delta\mathcal{Q}(s)$. The variation of the covariance-matrix elements with respect to the pulse amplitude $\Omega(t)$ and phase $\phi(t)$ can be expressed as a linear combination of the variation of these second-order moments with respect to the driving amplitude $\Omega(t)$ and phase $\phi(t)$,
@@
It can be seen from Eq. (18) that, to obtain the result in Eq. (17), we need to further calculate the variations of these second-order moments with respect to $\Omega(t)$ and $\phi(t)$. This can be achieved by taking the variation with respect to $\Omega(s)$ and $\phi(s)$ on both sides of Eq. (8), namely,
@@
Since $\mathcal{Q}(s)$ is a function of $s$, $\delta\dot{\mathbf{X}}(t)/\delta\mathcal{Q}(s)$ in Eq. (19) can be expressed as $\frac{d}{dt}\frac{\delta{\mathbf{X}}(t)}{\delta\mathcal{Q}(s)}$ Greiner1996FQ . Under the initial condition $\delta\mathbf{X}(0)/\delta\mathcal{Q}(s)=[0,...,0]_{10\times 1}^{\text{T}}$, the solution of Eq. (19) at the target time $T$ can be obtained as
@@
To know the explicit expressions for $\delta\mathbf{M}(s)/\delta\mathcal{Q}(s)$ and $\delta\mathbf{N}(s)/\delta\mathcal{Q}(s)$ in Eq. (A), we need to obtain the values of $\delta\alpha(s)/\delta\mathcal{Q}(s)$, $\delta\beta(s)/\delta\mathcal{Q}(s)$, and their Hermitian conjugate. Taking the variation with respect to $\mathcal{Q}(s)$ on both sides of Eqs. (6a) and (6b) as well as their Hermitian conjugate equations, we have
@@
$\mathbf{A}(t)=\left(\frac{\delta\alpha(t)}{\delta\mathcal{Q}(s)},\frac{\delta\beta(t)}{\delta\mathcal{Q}(s)},\frac{\delta\alpha^{\ast}(t)}{\delta\mathcal{Q}(s)},\frac{\delta\beta^{\ast}(t)}{\delta\mathcal{Q}(s)}\right)^{\text{T}},$ (22)
@@
$\left(\frac{\delta\alpha(s)}{\delta\Omega(s)},\frac{\delta\beta(s)}{\delta\Omega(s)},\frac{\delta\alpha^{\ast}(s)}{\delta\Omega(s)},\frac{\delta\beta^{\ast}(s)}{\delta\Omega(s)}\right)^{\text{T}}=\left(\frac{1}{2}ie^{-i\phi(s)},0,-\frac{1}{2}ie^{i\phi(s)},0\right)^{\text{T}},$ (26a)
$\left(\frac{\delta\alpha(s)}{\delta\phi(s)},\frac{\delta\beta(s)}{\delta\phi(s)},\frac{\delta\alpha^{\ast}(s)}{\delta\phi(s)},\frac{\delta\beta^{\ast}(s)}{\delta\phi(s)}\right)^{\text{T}}=\left(\frac{1}{2}\Omega(s)e^{-i\phi(s)},0,\frac{1}{2}\Omega(s)e^{i\phi(s)},0\right)^{\text{T}}.$ (26b)
Thus, we obtain the variation of the displacement amplitudes $\alpha(s)$ and $\beta(s)$ with respect to the driving amplitude $\Omega(s)$ and phase $\phi(s)$, respectively. Based on the value of $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}(s)$, we can further perform the gradient-descent algorithm until a satisfactory value of the loss function $\Delta X_{b}^{2}(\theta,T)$ is obtained.